Status Report

ISS On-Orbit Status 11 Feb 2002

By SpaceRef Editor
February 11, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below. Week 9 is underway for Expedition Four (“Week 10” in yesterday’s status was an error).

Before breakfast, all three crewmembers completed the Russian MedOps MO-10 “Hematocrit” examination, which analyzes blood samples for their hematocrit value (percentile ratio of volume of red blood cells to volume of whole blood). This test is performed every 60 days.

Major task assignment for today was review and familiarization with timeline and task details of the upcoming U.S. Airlock EVA, scheduled for Wednesday next week (2/20). A readiness review with ground specialists and crew is planned for this Friday, 2/15. After Bursch and Walz reviewed the tasks to be performed during the 6-hr. excursion, Onufrienko joined in for a joint walk-through of the timeline. The two flight engineers then completed about two hoursâ worth of OBT (on-board training) for SAFER (simplified aid for EVA rescue) operations, using the computerized joystick-operated “flight simulator”. They concluded the activities with a checkout of their SAFER units.

CDR Onufrienko installed the Russian GFI-10 “Molniya” payload, which had come up with Progress 6P. The experiment was mounted at SM window #1 for the geophysical study of atmosphere, ionosphere and magnetosphere electromagnetic interaction related to storms and seismic activities, using video-photometric systems (VFS-3M). Installation requires precise alignment [i.e., horizontal axis of payload parallel to longitudinal axis of SM, and line-of-sight axis 27 degrees below the local horizon]. After its porthole installation, the equipment operated automatically for several hours, controlled from the SM on-board equipment control system (SUBA). It was then deactivated, taken apart and stowed away. The GFI-10 experiment is conducted periodically in a total of six sessions during the current increment.

Yuri also performed systems service activities, e.g., removing and replacing dust filters in the SOTR thermal control system (two in the DC-1, four in the SM), inspection of autonomous Lab payloads, BRPK-1 condensate separator, SRVK condensate water processor and SVO water supply subsystem, and maintenance of SOSH life support subsystems.

FE-2 Bursch transferred data from his microbial colony growth analysis to the MEC (medical equipment computer). Usually, such data are obtained after 2 and 5 days of incubation, but the ground determined that the T+5 d samples were not required this time.

As reported on 2/8, the #2 ATU (audio terminal unit) in the Lab tripped its RPCM (remote power controller module) over the weekend of 2/2-2/3, akin to “blowing a fuse”. This has now evolved into a larger problem on board. Troubleshooting of the Lab RPCM5’s tripping by crew and ground over the past weekend shows that that there appears to be a larger malfunction in the RPCM, not in the ATU. MCC-H is gathering data and putting together a plan for tomorrow. Since removal and replacement (R&R) of the RPCM, which controls one of the LA MDMs (Lab systems computers), would require half of the Lab to be powered down, necessary steps are being considered with great circumspection. Among the systems down due to the loss of the LA1 MDM are the GNC1 MDM (guidance, navigation & control computer #1). The Ku-band was also temporarily offline but is now back up again. There are no actions for the crew for the immediate future. [The ATUs are major input/output units for the internal audio subsystem (IAS). There are two of them in the Lab module, one on each end, and ATU #1 continues to work. The IAS is a multichannel, multi-access, full-duplex (“both ways”) audio intercom network which allows the crewmembers to communicate with each other throughout the ISS. Its configuration is controlled by commands originating from an ATU panel, the PCS laptops, or the ground. Most importantly, caution and warning (C&W) tones are annunciated over the IAS to each ATU, and there is also a paging function available on the ATU].

All three crew performed their regular 2-hr. physical exercise on TVIS, RED and VELO cycle ergometer.

Today’s targets for the CEO imaging program were Irrawaddy River Delta (pass was just to the NW of this delta complex. Of interest: documenting the water color and extent of sediment plumes in this dry season and low water flow period), Bombay, India (a near nadir pass with good light and weather for mapping the details of this megacity and seaport on west coast of India), Angolan Biomass Burning (lower Namibia has remained dry for several weeks now making it possible for burning to begin. Of interest: possible smoke plumes to left of track after crossing the coast), Rift Triple Junction, Ethiopia (pass was just to the SW of this target. As ISS approached the Gulf of Aden from the SW, crew was to look just to the left track and map the complex of faults and folds found where the African and Arabian Plates meet), Lake Nasser, Toshka Lakes, (as the station approached the Nile valley from the SW, crew was advised to look slightly to the right of track and carefully map the large man-made Toshka Lake system spilling out into the desert west of Lake Nasser), E. Mediterranean Dust and Smog (of interest: during approach to southern Greece from the SW, view to the right of track over the Aegean Sea should show smog palls forming there), European Smog (at the coast of northern Italy, crew was to look to the right of track for smog accumulation over the basin of the Adriatic Sea), Eastern United States (of interest: oblique and limb views to the left of track of smog plumes forming and lingering over the Piedmont Plateau), and Tuamotu Archipelago (of interest: use of the ESC [electronic still camera] to map the numerous reefs and atolls here as ISS tracked northeastward across the heart of this extensive archipelago).

U.S. and Russian Segment Status (as of 2:24 pm EST today):

Environmental Control and Life Support (ECLSS):

  • Elektron O2 generator is On (16 Amp mode). Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is Off.
  • BMP Harmful Impurities unit: Both absorbent beds (Filters #1 & #2) in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 748, temperature (deg C) — 26.2, ppO2 (mmHg) — 158.2, ppCO2 (mmHg) — 2.56(?)
  • SM Transfer Compartment: Pressure (mmHg) — 743, temperature (deg C) — 20.5; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • FGB Cabin: Pressure (mmHg) — 744, temperature (deg C) — 21.3; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Node: Pressure (mmHg) — 733.07, temperature (deg C) — 22.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 735.22, temperature (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock: Pressure (mmHg) — Ê734.92, temperature (deg C) — n/a

(Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Directed position (125 degrees).
  • SM battery #6 is cycling; all other ( 7) SM batteries are in “partial charge” mode.
  • FGB battery #5 is offline; all other (5) batteries are in “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is Off.
  • LA1 MDM is Off; LA2 and LA3 MDMs are operating.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally, and Ku-band is operating nominally in open loop pointing mode.
  • Audio subsystem operating nominally (except for ATU #2; see above).
  • Video subsystem operating nominally (but VTR2 is non-functional).
  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port stow position (on Keep Alive power on both strings).
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 4:06 am EST):

  • Mean altitude — 384.99 km
  • Apogee — 388.9 km
  • Perigee — 381.0 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0005855
  • Orbits per 24-hr. day — 15.61
  • Decay rate — 480 m (mean) in last 24 hours
  • Solar Beta Angle: -8.2 deg (magnitude increasing)
  • Revolutions since FGB/Zarya launch (Nov. â98) — 18440
  • Current Flight Attitude — LVLH +XVV ZLV (local vertical/local horizontal: +X-axis in velocity vector; Z-axis in local vertical), with TEA (torque equilibrium attitude [pitch: -10 deg, yaw: -7 deg, roll: 0 deg]). Until 2/14.

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.